OpenStudy (anonymous):

-17b9 _________ 17b9 negative 17 to the power of 9 divided by 17 to the power of 9 ?

OpenStudy (mathstudent55):

\( \dfrac{-17b^9}{17b^9} \) \(= \dfrac{-17}{17} \times \dfrac{b^9}{b^9} \) \(= -1 \times 1\) \(= -1\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-30b5 _________ -30

OpenStudy (anonymous):

i need help on this one too

OpenStudy (mathstudent55):

\( \dfrac{-30b^5}{-30} \) \(= \dfrac{-30}{-30} \times \dfrac{b^5}{1} \) \(= 1 \times b^5 \) = \(b^5\)